### v1

# GUTs - how common are they?

Abstract
The individual fermion generations of the Standard Model fit neatly into a representation of a simple Grand Unified Theory gauge algebra. If Grand Unification is not realized in nature, this would appear to be a coincidence. We attempt to quantify how frequent this coincidence is among theories with group structure and fermion content similar to the Standard Model. We find that $\mathcal{O}(1/100)$ of consistent fermion representations similar to the Standard Model are unifiable, and discuss how the result depends on our definitions. This purely group-theoretical analysis may be taken as a bottom-up indication for Grand Unification, conceptually similar to a naturalness argument.
The Standard Model (SM) of particle physics unifies electromagnetic and weak forces into a single framework [1, 2, 3]. The SM gauge forces, in turn, can be unified into a more symmetric Grand Unified Theory (GUT) [4, 5, 6, 7, 8]. Intriguingly, the SM fermions fit neatly into $SU(5)$ representations [5] and if a right-handed neutrino is added, one generation of fermions fits exactly into the $\mathbf{16}$ representation of $SO(10)$ [8, 7]. This perfect fit seems to be too good to be a mere coincidence and is part of the appeal of GUTs. However, since symmetries and unification are driving concepts in physics, successfully constructing a GUT may be more of a result of our own preoccupations than an observation about nature.
In this letter, we try to quantify how surprised we should be at the ‘unifiability’ of the SM fermions. We construct a base set of theories that look similar to the SM and check what fraction of them can be embedded in a GUT. To obtain an answer, we need to define what we mean by ‘SM-like’ theories and which theories we consider to be ‘unifiable’. The result will depend on these arbitrary choices, but in a systematic way, allowing us to draw conservative conclusions.
Unifiability
We use a UV-agnostic, bottom-up approach for unifiability where we ask if a given set of observed fermions by itself is unifiable into representations of a simple GUT algebra, without the need for additional, hitherto unobserved fermions. The condition of no additional fermions provides closure to the problem and resembles the situation in the SM, where the known fermions unify into a representation of $SU(5)$. Also note that we do not consider gauge coupling unification, as it is only suggestive in the SM and depends on the scalar sector as well. Our group-theoretic definition of unification is therefore only a necessary condition, such that our results will be conservative in the sense that the fraction of actually unifying theories will be smaller.
Figure 1: Fraction $r$ of SM-like, (fully) chiral, anomaly-free fermion representations that are unifiable into a representation of a simple GUT gauge algebra, as a function of the maximal considered fermion dimension $D_{\mathrm{max}}$. Representations are restricted to $U(1)$ charges of $|Q|<10$ and at most $\tilde{S}=4$ identical irreducible representations under the semi-simple part of the algebra.
Standard Model-like theories
To assess the rarity of the unifiability property of the SM, we construct a set of theories that includes and generalizes the SM. We consider the essential observational facts of the SM to be i) three gauge forces corresponding to a reductive gauge algebra with a rank-3 semi-simple part, ii) $D=15$ fermions per generation, iii) each generation is itself anomaly free, iv) the fermions carry integer hypercharges $|Q|\leq 6$, v) the fermion representation is chiral.
Based on these observations, we consider self-consistent (ie. anomaly-free), chiral  [^1] We call a set of fermions ‘chiral’ if it is a complex representation of the gauge algebra, and ‘fully chiral’ if it contains no real (vector-like) subset. representations of the SM gauge algebra, with dimension $D\leq D_{\mathrm{max}}$ and integer charges $|Q|\leq 10$ as base set of SM-like theories. We also restrict the number $\tilde{S}$ of identical irreducible representations under the semi-simple part of the gauge algebra to $\tilde{S}\leq 4$. The result depends on these arbitrary choices, which we will discuss before also considering different gauge algebras.
Likelihood of unifiability
We quantify the likelihood that a theory with an anomaly-free representation of dimension $D$ unifies into a simple GUT by the ratio of unifiable representations over all anomaly-free representations up to dimension $D_{\rm max}\geq D$. This is shown in Figure 1 for the SM gauge algebra. The dependence of this likelihood on $D_{\rm max}$ will be discussed with our results.
I Methods
There are two steps to assessing how common unifiability is among SM-like theories. In a bottom up approach, we first construct the set of all consistent SM-like theories. Then, we check unifiability for each of them, using the SuperFlocci [9] code. Using the GroupMath [10] code, we can verify and extend our results by a top-down determination of branchings of all candidate GUTs.
Constructing anomaly-free representations
The construction of the set of all anomaly-free representations is performed in three steps:
  1. i)
    Find all anomaly-free representations of the semi-simple part of the gauge algebra ($SU(3)\times SU(2)$ in the SM) up to dimension $D_{\rm max}$.
  2. ii)
    Assign integer $U(1)$ charges within a predefined range $|Q|\leq Q_{\rm max}$ to all representations under the semi-simple part of the algebra and keep those that satisfy the anomaly cancellation conditions (gravitational and gauge anomalies). [^2] Note that since we work at the level of Lie algebras we do not check for global anomalies, such as the anomaly associated with an odd number of fermion doublets charged under $SU(2)$ in four dimensions [11]. Such global anomalies depend on the global structure of the Lie group and cannot be determined from the Lie algebra alone as there is no one-to-one correspondence between Lie groups and algebras. However, we have checked that excluding representations with an odd number of $SU(2)$ doublets in the base set changes our results by an $\mathcal{O}(1)$ number (the results in eqns. (1) and (2) change to $37/106763$ and $1/37$, respectively).
  3. iii)
    Filter out equivalent representations. We consider representations equivalent if they differ only by an integer rescaling of the $U(1)$ charge or are conjugate representations of each other (or a combination of both). For this reason we keep only representations where the greatest common divisor of all charges is one. An example of equivalent $(SU(3),SU(2))_{U(1)}$ representations is
    $\begin{split}&(\mathbf{3},\mathbf{2})_{0}\oplus(\mathbf{\bar{3}},\mathbf{1})_{-1}\oplus(\mathbf{\bar{3}},\mathbf{1})_{1}\\ &\qquad\qquad\quad\quad\sim(\mathbf{\bar{3}},\mathbf{2})_{0}\oplus(\mathbf{3},\mathbf{1})_{-2}\oplus(\mathbf{3},\mathbf{1})_{2}\,,\end{split}$
    since they are conjugate representations with rescaled $U(1)$ charges.
While i) and iii) are easily implemented with Mathematica packages such as SuperFlocci [9] or GroupMath [10], ii) is a challenging combinatoric problem since the number of possible charge assignments grows exponentially with the number of fermions. In order to deal with this large number of possible charge assignments we use compiled Mathematica code and simplify the problem for a given semi-simple representation in the following way. We generate charge assignments for blocks of identical irreducible representations within a given candidate representation in a lexicographic order that does not go through permutations of charges within one block. Additionally we split the candidate representation in two and compute anomaly coefficients for charge assignments in each part separately. Finally we match those assignments which add up to zero when combined from the two halves. Despite the above simplifications we cannot handle semi-simple representations which are composed of a large number of irreducible representations. For this reason we restrict the base set to representations that contain no more than $\tilde{S}\leq 4$ equal semi-simple irreducible representations. This is on the one hand necessary to limit the number of $U(1)$ charge assignments when extending the analysis to large fermion dimensions and on the other hand mimics the situation in the SM where no semi-simple irreducible representation appears more than twice in a generation. We further demand that all forces have at least one particle charged under them.
SuperFlocci [9] – bottom-up determination of unifiability
SuperFlocci is a Mathematica package that “takes any reductive gauge algebra and fully-reducible fermion representation, and outputs all semi-simple gauge extensions under the condition that they have no additional fermions, and are free from local anomalies” [9]. A theory is unifiable according to the definition laid out in the introduction exactly if SuperFlocci finds a simple gauge extension.
Figure 2: Number of consistent fermion representations up to dimension $D_{\mathrm{max}}$. Dashed lines correspond to consistent representations, while solid lines only count those that unify. The green and red colors restrict the analysis to chiral or fully chiral fermion representations, respectively. We also show some results without chirality restriction in blue. Left: Results for the Standard Model gauge algebra. The jumps in the number of unifiable theories occur when unification into $SO(10)$ ($SU(6)$ and $E_{6}$) become possible at $D_{\rm max}=16$ ($D_{\rm max}=27$). Right: Results for $\mathrm{SU}(3)\times\mathrm{U}(1)$.
GroupMath [10] – top-down construction of unifiable representations
As a second approach, we consider all candidate unified (ie. simple) gauge algebras that have (non-singlet) representations with dimension $\leq D_{\mathrm{max}}$. We use the GroupMath Mathematica package to find all decompositions of all candidate GUT representations to the SM gauge algebra. We apply the same filters to the SM representations as to the base set and assign $U(1)$-charges according to the rules determined by GroupMath. This is an independent top-down check on the number of unifiable theories and gives the same result as the bottom-up analysis using SuperFlocci. It is, however, more efficient and allows to extend the analysis to larger fermion dimensions.
Examples
To illustrate our approach we provide a few examples for consistent theories that we do or do not consider in the base set of SM-like theories, and that are or are not unifiable.
The smallest fermion representation free from local anomalies under the SM gauge algebra $\mathrm{SU}(3)\times\mathrm{SU}(2)\times\mathrm{U}(1)$, which has particles charged under all three forces, is:
$(\mathbf{1},\mathbf{2})_{0}\oplus(\mathbf{3},\mathbf{1})_{-1}\oplus(\bar{\mathbf{3}},\mathbf{1})_{1}\,.$
It unifies into an $\mathbf{8}$ of $\mathrm{Sp}(8)$, but is not chiral. [^3] Even though this representation is not complex (and hence not counted as chiral here), a direct mass term for the $(\mathbf{1},\mathbf{2})_{0}$ is not allowed. The first chiral representation appears at $D=12$:
$(\mathbf{3},\mathbf{2})_{0}\oplus(\bar{\mathbf{3}},\mathbf{1})_{-1}\oplus(\bar{\mathbf{3}},\mathbf{1})_{1}\,.$
It is not unifiable. The smallest chiral, unifiable representation of the SM gauge algebra is the single-generation SM at $D=15$:
$(\mathbf{1},\mathbf{1})_{-6}\oplus(\mathbf{1},\mathbf{2})_{3}\oplus(\bar{\mathbf{3}},\mathbf{2})_{-1}\oplus(\mathbf{3},\mathbf{1})_{-2}\oplus(\mathbf{3},\mathbf{1})_{4}\,.$
II Results
Fig. 1 shows the ratio $r$ of the number of unifiable representations over the number of all SM-like fermion representations of the SM gauge algebra, depending on the fermion dimension $D\leq D_{\mathrm{max}}$ and restricted to representations with at most $\tilde{S}\leq 4$ identical irreducible representations under the semi-simple part of the algebra and $U(1)$ charges $|Q|\leq 10$. The lowest dimensional unifiable representation appears at $D=15$, and roughly $10^{-3}$ of all consistent chiral representations with $D\leq 15$ are unifiable. This ratio drops as $D_{\mathrm{max}}$ is raised, as the number of consistent theories grows faster than the number of unifiable ones, as can be seen in the left panel of Figure 2.
Figure 3: Dependence of the unifiable fraction on arbitrary definitions of SM-likeness. Left: Cut on maximal considered integer charge $Q_{\rm max}$. Right: Cut on the number $\tilde{S}$ of equal irreducible representations of the semi-simple part of the gauge algebra.
Picking, for concreteness, $D_{\mathrm{max}}=20$ as arbitrary cut on what we consider to be SM-like, we find
[…]
This number depends sensitively on the arbitrary choices of $D_{\mathrm{max}}$ and $Q_{\mathrm{max}}$. We also find that if we do not restrict $\tilde{S}$, the base set is inflated by theories with a large number of semi-simple singlets or $SU(2)$ doublets for larger $D_{\rm max}$. Figs. 1 and 3 show how the result depends on these arbitrary cuts. For all of $Q_{\mathrm{max}}$, $D_{\mathrm{max}}$ and $\tilde{S}$, there is a clear trend of falling unifiable fraction for more general definitions of SM-like. As can be seen in both panels of Fig. 3, more restrictive cuts lead to more conservative estimates, i.e. larger values of $r$. The most conservative result is obtained when imposing the tightest restrictions that still include the SM, $D\leq 15,|Q|\leq 6,\tilde{S}\leq 2$:
[…]
Instead of considering chiral, ie. not completely vector-like (VL), fermion representations, we can restrict the analysis to fully chiral representations (without any singlets or VL particles). The resulting unifiable fraction $r$ is also shown in Fig. 1 and paints a different picture: At $D_{\mathrm{max}}\sim D_{\mathrm{SM}}$, almost all fully chiral anomaly-free fermion representations are unifiable, while for $D_{\mathrm{max}}=20$, we find $r\sim 10^{-2}$, similar to the most conservative result for chiral theories. Note that the single-generation SM is the smallest unifiable, fully chiral fermion representation of the SM gauge algebra.
Relaxing all chirality restrictions is not very informative, as both the base set and the unifiable set are then dominated by the large number of completely VL representations (see blue curve in the left panel of Fig. 2), which are very unlike the SM. The number of theories quickly becomes difficult to handle computationally, but for $D_{\mathrm{max}}\leq 20$ we find $r\sim 1$.
III Discussion
To check whether the SM gauge algebra itself is special when it comes to unifiability, we repeated the analysis for different semi-simple parts of the gauge algebra. In the case of $SU(3)\times U(1)$, we need to go to $D\geq 27$ to find the first non-VL, unifiable representation, where we find a unifiable fraction of order $10^{-5}$, similar to the SM result, even if we allow for $\tilde{S}=7$ in order not to exclude any unifiable representations (see the right panel of Fig. 2). For the remaining rank-2 gauge algebras $\{SO(5),SU(2)\times SU(2),SP(4),G_{2}\}\times U(1)$ we do not find any chiral, unifiable fermion representations with $D\leq 30$. The same is true for rank-3 algebras, such as $SU(4)\times U(1)$, with the exception of the SM gauge algebra. This is partially due to the increasing dimension of the smallest representations of these algebras. At the same time the base set of chiral theories is still growing exponentially since $U(1)$ charge assignments can be used to make sets of five or more fermions chiral [12, 13]. Thus, the number of chiral, unifiable representations is typically smaller than for the SM gauge algebra. Any result for the SM gauge algebra can therefore be considered conservative, since allowing for different gauge algebras would reduce the fraction of chiral, unifiable theories.
Let us also mention that different definitions of unifiability are possible. Our approach of demanding that the fermions fit neatly into a representation of the GUT that includes no further fermions may be relaxed. From a practical perspective, there is no harm if a GUT predicts hitherto unobserved VL fermions – those may be heavy. This is the case for instance for the right handed neutrinos predicted in $SO(10)$ grand unification [7]. A new condition to provide closure to the problem would be needed (eg. an arbitrary cut on the number of inferred VL fermions). Using our existing results, we can estimate the impact of relaxing the unifiability criterion by considering all theories as unifiable that unify when adding VL fermions up to a total fermion dimension of $30$: In this case, the unifiable fraction of Fig. 1 increases by a factor of up to (2) 4 in the (fully) chiral case, and the first chiral unifiable theory appears at $D=12$. We leave the study of stronger unification criteria (eg. unification of fermions into an irreducible representation, as in $SO(10)$ GUTs) to future work.
It is possible that the unifiability of the SM is not a consequence of originating from a grand unified theory but a consequence of another property that correlates with unifiability. As seen in Fig. 1, being fully chiral is one such property that is highly correlated with unifiability in the immediate neighborhood of the SM. Ironically, when adding our top-down expectations that only the fully chiral fermions stay light, the unifiability of the observed fermions is no longer by itself a strong indication for grand unification.
IV Conclusion
In this letter, we consider the single-generation SM as one among many similar consistent theories. From this starting point, the observation that the SM fermion representation is unifiable may seem surprising. Here we try to find a quantitative answer to the question of how surprised we should actually be. We find that the unifiability of fermions into a representation of a simple unified algebra, as it occurs in the SM into a $SU(5)$ GUT, is rare among SM-like chiral theories at the 1/100 level. Stronger statements are possible when allowing for larger fermion dimension, charges or number of distinct irreducible representations than the SM. On the other hand, when restricting the analysis to only include fully chiral theories, unifiability is common among anomaly-free fermion representations not larger than the SM. The argument presented here can be taken as a purely group-theoretical indication for Grand Unification, conceptually similar to a naturalness argument. However, the absence of a probability measure in the space of theories hampers a probabilistic interpretation of our results.
Acknowledgements.
We thank Joseph Tooby-Smith for collaboration in the early stages of this project and coding support throughout. We also thank Andreas Trautner and Joseph Tooby-Smith for valuable comments on the manuscript. MR is supported by the NSF grant PHY-2014071. The work of MR was performed in part at the Aspen Center for Physics, which is supported by a grant from the Simons Foundation (1161654, Troyer).